# Reductions in local certification

### v1

@@ Abstract.
Local certification is a topic originating from distributed computing, where a prover tries to convince the vertices of a graph $G$ that $G$ satisfies some property $\mathcal{P}$. To convince the vertices, the prover gives a small piece of information, called certificate, to each vertex, and the vertices then decide whether the property $\mathcal{P}$ is satisfied by just looking at their certificate and the certificates of their neighbors. When studying a property $\mathcal{P}$ in the perspective of local certification, the aim is to find the optimal size of the certificates needed to certify $\mathcal{P}$, which can be viewed a measure of the local complexity of $\mathcal{P}$.
@@
The authors are partially supported by the French ANR Project TWIN-WIDTH (ANR-21-CE48-0014-01), and by LabEx PERSYVAL-lab (ANR-11-LABX-0025).
@@ 1. Introduction
If the graph indeed satisfies $\mathcal{P}$, all vertices must accept the instance. If the graph does not satisfy $\mathcal{P}$, then for any possible assignment of certificates, at least one vertex must reject the instance. For $n\in\mathbb{N}$, the local complexity of $\mathcal{P}$ is the minimum $f(n)$ such that for every $n$-vertex graph $G$, the property $\mathcal{P}$ can be locally certified in $G$ with certificates of at most $f(n)$ bits per vertex. We can also study more generally the local complexity of a property $\mathcal{P}$ within a graph class $\mathcal{C}$, by which we mean that the minimum above is taken over all $n$-vertex graphs from $\mathcal{C}$.
@@
  • $\Theta(\log n)$: this includes non-bipartiteness, acyclicity; planarity [9] and more generally bounded genus [10, 6];
@@
Many of the results mentioned above (either lower bounds or upper bounds) are proved using rather ad-hoc arguments that do not seem to be applicable to a large number of graph properties at once. On the other hand, a number of positive results (upper bounds on the local complexity) have been obtained via so-called “meta-theorems”, that is theorems applicable to a wide range of problems and graph classes. It was proved in [12] that any graph class of bounded treewidth which is expressible in monadic second order logic has local complexity $O(\log^{2}n)$ (this implies in particular that for any fixed $k$, the class of graphs of treewidth at most $k$ has local complexity $O(\log^{2}n)$). Similar meta-theorems involving graph properties expressible in first-order logic were proved for graphs of bounded treedepth [8] and graphs of bounded cliquewidth [11]. Another type of meta-theorem was obtained in [17]: any property that can be decided in linear time has an interactive proof labeling scheme with compact certificates (in this setting several rounds of interactions are allowed between the prover and the vertices).
Distributed algorithms in graphs are considered to be efficient if they run in a polylogarithmic number of rounds. Therefore, the class of all properties whose local complexity is polylogarithmic in the size of the graph can be seen as a distributed analogue of the class NP. However, a major feature of NP has so far been missing in the toolbox of local certification: the possibility of doing hardness reductions. The purpose of this paper is to introduce a framework allowing to use classical NP-hardness reductions to prove automatic lower bounds on the local complexity of graph properties. Our techniques apply whenever the reduction is sufficiently local. In particular, using the result of Göös and Suomela [14] stating that non-3-colorability has local complexity $\Omega(n^{2}/\log n)$, we prove that a number of classical coNP-hard problems have polynomial local complexity. The table in Figure 1 sums up our lower bounds.
@@
Domatic number at most $2$ Graphs with $\Delta\leqslant 8$ $\Omega\left(\frac{n}{\log n}\right)$ 5.4
Non-existence of a cubic subgraph Graphs with $\Delta\leqslant 7$ $\Omega\left(\frac{n}{\log n}\right)$ 5.5
Non-existence of a partition in $k$ acyclic subgraphs ($k\geqslant 3$) General graphs $\Omega\left(\frac{n}{\log n}\right)$ 5.9
Existence of a monochromatic triangle in every $2$-edge-coloring Graphs with $\Delta\leqslant 40$ $\Omega\left(\frac{n}{\log n}\right)$ 5.10
@@
Figure 1. Summary of our lower bounds. In this table, $\Delta$ denotes the maximum degree.
Note that most of the lower bounds in the table of Figure 1 hold in a bounded-degree graph class (so in general graphs too). However, in a bounded-degree graph class, every property has local complexity $O(n\log n)$. Indeed, $O(n\log n)$ bits are sufficient to encode all the edges of a bounded-degree graph (using adjacency lists), and we can just write the description of the whole graph in the certificate of every vertex. Thus, most of the results above in bounded-degree classes are optimal (within a polylogarithmic factor). The two exceptions are the local complexity of non-3-colorability in planar graphs of maximum degree 4, and the local complexity of non-hamiltonicity in graphs of maximum degree 4. We believe that in both cases the local complexity should be linear. For the non-existence of a partition into $k$ acyclic subgraphs, we only have a quasi-linear lower bound, which holds in general graphs. This is probably also not optimal (because in general graphs the universal upper bound is $O(n^{2})$). Finally, note also that, except for $k$-colorability, our quasi-linear lower bounds that holds in bounded-degree classes remain quasi-linear in general graphs. Again, this is probably not optimal: an interesting open problem would be to determine whether each of them can be improved to quadratic in general graphs.
@@
In Section 7, we present a small variant of our general reduction theorem (namely, we give a relaxation of one of the conditions needed to apply it, that holds for properties on bounded-degree graph classes). We present again a direct application. Finally, in Section 8, we show that there is no hope for a much more general theorem stating that any coNP-hard problem has polynomial local comlexity: more precisely, we give an example of a coNP-hard problem that has only logarithmic local complexity. We also show that our reduction framework applies not only to coNP-hard problems, by giving examples of reductions showing polynomial lower bounds on the local complexity of some properties that are in P.
@@ 2.1. Graphs
In this paper, graphs are assumed to be simple, loopless, undirected, and connected. The distance between two vertices $u$ and $v$ in a graph $G$, denoted by $d_{G}(u,v)$ is the minimum length of a path between $u$ and $v$ (that is, its number of edges). The neighborhood of a vertex $v$ in a graph $G$, denoted by $N_{G}(v)$ (or $N(v)$ is $G$ if clear from the context), is the set of vertices at distance exactly 1 from $v$. The closed neighborhood of $v$, denoted by $N_{G}[v]:=\{v\}\cup N_{G}(v)$, is the set of vertices at distance at most 1 from $v$. For a set $S$ of vertices of $G$, we define $N_{G}[S]:=\bigcup_{v\in S}N_{G}[v]$. Finally, for every $k\geqslant 1$, we define $N_{G}^{k}[v]$ as the set of vertices at distance at most $k$ from $v$ (in particular, we have $N_{G}^{1}[v]=N_{G}[v]$).
@@ 2.2. Local certification
A verifier $\mathcal{A}$ is a function that takes a graph $G$, a proof $P$ for $G$, and a vertex $v\in V(G)$ as inputs, and outputs an element of $\{0,1\}$. We say that $v$ accepts the instance if $\mathcal{A}(G,P,v)=1$ and that $v$ rejects the instance if $\mathcal{A}(G,P,v)=0$.
@@ 3. Non-3-colorability
In what follows, we will use the following result, proved by Göös and Suomela in [14]:
@@
Since we will use several variants of this result, we give a short description of their construction. For any integer $k$, and any two subsets $A\subseteq[2^{k}]^{2}$ and $B\subseteq[2^{k}]^{2}$, Göös and Suomela construct a graph $G_{A,B}$ on $n=\Theta(2^{k})$ vertices such that:
@@
    Each edge between $V_{A}$ and $V_{B}$ in $G_{A,B}$ connects a vertex of $S_{A}$ to a vertex of $S_{B}$, there are at most $O(k)$ such edges, and the subgraph of $G_{A,B}$ induced by the special vertices $S_{A}\cup S_{B}$ is independent of the choice of $A$ and $B$.
@@
Given labelings of two graphs $G_{A,B}$ and $G_{C,D}$, we say that the labelings agree on their special vertices, if for every $1\leqslant i\leqslant s$ the label of $a_{i}$ in $G_{A,B}$ coincides with the label of $a_{i}$ in $G_{C,D}$, and the label of $b_{i}$ in $G_{A,B}$ coincides with the label of $b_{i}$ in $G_{C,D}$.
Göös and Suomela [14] observed that since there are only $O(\log n)$ special vertices, if non-3-colorability can be certified locally with certificates of size $o(n^{2}/\log n)$, the total number of bits of certificates assigned to the special vertices is $o(n^{2})$. Since there are $2^{(2^{k})^{2}}=2^{\Omega(n^{2})}$ choices for the sets $A$, it follows from the pigeonhole principle that there exist two distinct sets $A,B$ such that the certificates of $G_{A,\bar{A}}$ and $G_{B,\bar{B}}$ agree on their special vertices. As $G_{A,\bar{A}}$ and $G_{B,\bar{B}}$ are both non-3-colorable, in any local certification scheme for 3-non-colorability in these two graphs, each vertex accepts the instance upon reading its local view (i.e., the certificates of its closed neighborhood).
Note that at least one of $A\cap\bar{B}$ and $\bar{A}\cap B$ is non-empty, say $A\cap\bar{B}\neq\emptyset$ by symmetry. Consider the graph $G_{A,\bar{B}}$, and assign to the vertices of $V_{A}$ their certificates in $G_{A,\bar{A}}$, and to the vertices of $V_{\bar{B}}$ their certificate in $G_{B,\bar{B}}$. Observe that the local view of each vertex of $G_{A,\bar{B}}$ is the same as in $G_{A,\bar{A}}$ or $G_{B,\bar{B}}$, and thus each vertex accepts the instance. Since $G_{A,\bar{B}}$ is 3-colorable (because $A\cap\bar{B}\neq\emptyset$), we obtain a contradiction.
@@
For any integer $k$, and any two subsets $A\subseteq[2^{k}]^{2}$ and $B\subseteq[2^{k}]^{2}$, we consider the graph $G_{A,B}$ of Göös and Suomela [14] introduced above, and we let $H_{A,B}=f(G_{A,B})$. Note that $H_{A,B}$ contains $n=\Theta(2^{2k})$ vertices, and moreover:
@@
    Each edge between $V_{A}$ and $V_{B}$ in $G_{A,B}$, connects a vertex of $S_{A}$ to a vertex of $S_{B}$, and the subgraph of $G_{A,B}$ induced by the special vertices $S_{A}\cup S_{B}$ is independent of the choice of $A$ and $B$.
@@
The fact that the sets $S_{A}$ and $S_{B}$ still have size $O(\log n)$ in $H_{A,B}$ follows from item 3 in the properties of $G_{A,B}$ (the fact that there are only $O(\log n)$ edges with one endpoint in $S_{A}$ and one endpoint in $S_{B}$).
As above, if non-3-colorability can be certified locally with certificates of size $o(n/\log n)$ in graphs of maximum degree 4, the total number of bits of certificates assigned to the special vertices of $H_{A,B}$ is $o(n)$. Since there are $2^{(2^{k})^{2}}=2^{\Omega(n)}$ choices for the sets $A$, it follows from the pigeonhole principle that there exist two distinct sets $A,B$ such that the certificates of $G_{A,\bar{A}}$ and $G_{B,\bar{B}}$ agree on their special vertices. We then obtain a contradiction using the same fooling set argument as above. ∎
@@
We now consider $P^{\prime}_{A,B}=f(P_{A,B})$ (the graph obtained from $P_{A,B}$ by applying the maximum degree 4 reduction of the proof of Theorem 3.2). This graph is a planar graph with maximum degree 4, is 3-colorable if and only if $A\cap B\neq\emptyset$, has $n=O(2^{4k})$ vertices, but only contains $O(\log^{2}n)$ special vertices. There are $2^{(2^{k})^{2}}=2^{2^{2k}}=2^{\Omega(\sqrt{n})}$ choices for the set $A$, so if we can locally certify non-3-colorability with certificates of size $o(n^{1/2}/\log^{2}n)$ then the total number of bits of certificates assigned to the special vertices is $o(n^{1/2})$ and thus there are two distinct sets $A,B$ such that the certificates of $P^{\prime}_{A,\bar{A}}$ and $P^{\prime}_{B,\bar{B}}$ agree on their special vertices. We then obtain a contradiction using the same fooling set argument as above. ∎

### v2

@@ Abstract.
Local certification is a topic originating from distributed computing, where a prover tries to convince the vertices of a graph $G$ that $G$ satisfies some property $\mathcal{P}$. To convince the vertices, the prover gives a small piece of information, called certificate, to each vertex, and the vertices then decide whether the property $\mathcal{P}$ is satisfied by just looking at their certificate and the certificates of their neighbors. When studying a property $\mathcal{P}$ in the perspective of local certification, the aim is to find the optimal size of the certificates needed to certify $\mathcal{P}$, which can be viewed as a measure of the local complexity of $\mathcal{P}$.
@@
The authors are partially supported by the French ANR Project TWIN-WIDTH (ANR-21-CE48-0014-01), and ENEDISC (ANR-24-CE48-7768-01) by LabEx PERSYVAL-lab (ANR-11-LABX-0025). A preliminary version of this paper appeared in the proceedings of the 51st International Workshop on Graph-Theoretic Concepts in Computer Science (WG 2025).
@@ 1. Introduction
Note that there is no requirement on the computational power of the prover and the vertices. If the graph indeed satisfies $\mathcal{P}$, all vertices must accept the instance. If the graph does not satisfy $\mathcal{P}$, then for any possible assignment of certificates, at least one vertex must reject the instance. For $n\in\mathbb{N}$, the local complexity of $\mathcal{P}$ is the minimum $f(n)$ such that for every $n$-vertex graph $G$, the property $\mathcal{P}$ can be locally certified in $G$ with certificates of at most $f(n)$ bits per vertex. We can also study more generally the local complexity of a property $\mathcal{P}$ within a graph class $\mathcal{C}$, by which we mean that the minimum above is taken over all $n$-vertex graphs from $\mathcal{C}$.
@@
  • $\Theta(\log n)$: this includes non-bipartiteness, acyclicity, planarity [10] and more generally bounded genus [11, 7];
@@
Many of the results mentioned above (either lower bounds or upper bounds) are proved using rather ad-hoc arguments that do not seem to be applicable to a large number of graph properties at once. On the other hand, a number of positive results (upper bounds on the local complexity) have been obtained via so-called “meta-theorems”, that is theorems applicable to a wide range of problems and graph classes. It was proved that any graph class of bounded treewidth which is expressible in monadic second order logic has local complexity $O(\log^{2}n)$ [13]. This implies in particular that for any fixed $k$, the of treewidth at most $k$ has local complexity $O(\log^{2}n)$. Similar meta-theorems involving graph properties expressible in first-order logic were proved for graphs of bounded treedepth [9] and graphs of bounded cliquewidth [12]. Another type of meta-theorem was obtained in [19]: any property that can be decided in linear time has an interactive proof labeling scheme with compact certificates (in this setting several rounds of interactions are allowed between the prover and the vertices).
Proof labeling schemes in graphs are considered to be efficient if their complexity is polylogarithmic in the size of the graph. The class of all properties whose local complexity is polylogarithmic in the size of the graph can be seen as a distributed analogue of the class NP. However, a major feature of NP has so far been missing in the toolbox of local certification: the possibility of doing hardness reductions. The purpose of this paper is to introduce a framework allowing to use classical NP-hardness reductions to prove automatic lower bounds on the local complexity of graph properties. Our techniques apply whenever the reduction is sufficiently local. In particular, using the result of Göös and Suomela [15] stating that non-3-colorability has local complexity $\Omega(n^{2}/\log n)$, we prove that a number of classical coNP-hard problems have polynomial local complexity. Table 1 sums up our lower bounds.
@@
Domatic number at most $k$ ($k\geqslant 2$) Graphs with $\Delta\leqslant 5k^{2}-4k$ $\Omega\left(\frac{n}{\log n}\right)$ 5.4
Non-existence of a cubic subgraph Graphs with $\Delta\leqslant 7$ $\Omega\left(\frac{n}{\log n}\right)$ 5.6
Non-existence of a partition in $k$ acyclic subgraphs ($k\geqslant 3$) General graphs $\Omega\left(\frac{n}{\log n}\right)$ 5.10
Existence of a monochromatic triangle in every $2$-edge-coloring Graphs with $\Delta\leqslant 40$ $\Omega\left(\frac{n}{\log n}\right)$ 5.11
@@
Table 1. Summary of our lower bounds. In this table, $\Delta$ denotes the maximum degree.
Note that most of the lower bounds in Table 1 hold in a bounded-degree graph class (so in general graphs too). However, in a bounded-degree graph class, every property has local complexity $O(n\log n)$. Indeed, $O(n\log n)$ bits are sufficient to encode all the edges of a bounded-degree graph (using adjacency lists), and we can just write the description of the whole graph in the certificate of every vertex. Thus, most of the results above in bounded-degree classes are optimal (within a polylogarithmic factor). The two exceptions are the local complexity of non-3-colorability in planar graphs of maximum degree 4, and the local complexity of non-hamiltonicity in graphs of maximum degree 4. We believe that in both cases the local complexity should be linear. For the non-existence of a partition into $k$ acyclic subgraphs, we only have a quasi-linear lower bound, which holds in general graphs. This is probably also not optimal (because in general graphs the universal upper bound is $O(n^{2})$). Finally, note also that, except for $k$-colorability, our quasi-linear lower bounds which hold in bounded-degree graph classes remain quasi-linear in general graphs. Again, this is probably not optimal: an interesting open problem would be to determine whether each of them can be improved to quadratic in general graphs.
@@
In Section 7, we present a small variant of our general reduction theorem (namely, we give a relaxation of one of the conditions needed to apply it, that holds for properties on bounded-degree graph classes). We present again a direct application. Finally, in Section 8, we show that there is no hope for a much more general theorem stating that any coNP-hard problem has polynomial local complexity: more precisely, we give an example of a coNP-hard problem that has only logarithmic local complexity. We also show that our reduction framework applies not only to coNP-hard problems, by giving examples of reductions showing polynomial lower bounds on the local complexity of some properties that are in P.
@@ 2.1. Graphs
In this paper, graphs are assumed to be simple, loopless, undirected, and connected. The distance between two vertices $u$ and $v$ in a graph $G$, denoted by $d_{G}(u,v)$ is the minimum length of a path between $u$ and $v$ (that is, its number of edges). The neighborhood of a vertex $v$ in a graph $G$, denoted by $N_{G}(v)$ (or $N(v)$ if $G$ is clear from the context), is the set of vertices at distance exactly 1 from $v$. The closed neighborhood of $v$, denoted by $N_{G}[v]:=\{v\}\cup N_{G}(v)$, is the set of vertices at distance at most 1 from $v$. For a set $S$ of vertices of $G$, we define $N_{G}[S]:=\bigcup_{v\in S}N_{G}[v]$. Finally, for every $k\geqslant 1$, we define $N_{G}^{k}[v]$ as the set of vertices at distance at most $k$ from $v$ (in particular, we have $N_{G}^{1}[v]=N_{G}[v]$).
@@ 2.2. Local certification
A verifier $\mathcal{A}$ is a function that takes a graph $G$, a proof $P$ for $G$, and a vertex $v\in V(G)$ as inputs, and outputs an element of $\{0,1\}$. We say that $v$ accepts the instance if $\mathcal{A}(G,P,v)=1$ and that $v$ rejects the instance if $\mathcal{A}(G,P,v)=0$. In the definition we do not require that $\mathcal{A}$ is efficiently computable or even decidable (but in practice it will almost always be the case).
@@
A second difference between the setting of this paper and that of Göös and Suomela [15] is that vertices only look at distance at their neighbors 1 before accepting or rejecting the instance, while in [15] they can look at their neighborhood at distance $r$, for some constant $r$ called the local horizon. The lower bounds of Göös and Suomela hold in this more general setting, and so do the result in our paper (most of the time, without any modification to the proofs). However, there are quite a few places where having an arbitrary local horizon $r$ adds a layer of technicality on top of some already quite technical statements and proofs. So, for the sake of readability we have chosen to only state and prove our results in this paper in the case where the local horizon $r=1$, which already contains most of the difficulty of the problems we consider.
@@ 3. Non-3-colorability
In what follows, we will use the following result, proved by Göös and Suomela [15]:
@@
The full description of the construction in [15] takes over 6 pages (including several very helpful figures), so we do not describe it completely here. But since we will need several variants of this result, we give a short description of the main properties of the construction (the interested reader is referred to [15] for more details). Let $k$ be an integer, and let $I=\{1,\ldots,2^{k}\}$. For any two subsets $A\subseteq I\times I$ and $B\subseteq I\times I$, Göös and Suomela construct a graph $G_{A,B}$ on $n=\Theta(2^{k})$ vertices such that:
@@
    Each edge between $V_{A}$ and $V_{B}$ in $G_{A,B}$ connects a vertex of $S_{A}$ to a vertex of $S_{B}$, there are at most $O(k)$ such edges, and the subgraph of $G_{A,B}$ induced by the special vertices $S_{A}\cup S_{B}$ is independent of the choice of $A$ and $B$ (in particular $s$ does not depend on $A$ and $B$).
@@
Given assignments of certificates to the vertices of two graphs $G_{A,B}$ and $G_{C,D}$, we say that the certificates agree on their special vertices, if for every $1\leqslant i\leqslant s$ the certificate of $a_{i}$ in $G_{A,B}$ coincides with the certificate of $a_{i}$ in $G_{C,D}$, and the certificate of $b_{i}$ in $G_{A,B}$ coincides with the certificate of $b_{i}$ in $G_{C,D}$.
For completeness we now sketch the argument of Göös and Suomela [15] showing that non-3-colorability cannot be certified locally with certificates of size $o(n^{2}/\log n)$. Consider $G_{A,\bar{A}}$ for some set $A\subseteq I\times I$, where $\bar{A}$ denotes the complement of $A$. Note that (4) implies that $G_{A,\bar{A}}$ is not 3-colorable.
Since there are only $O(\log n)$ special vertices, if non-3-colorability can be certified locally with certificates of size $o(n^{2}/\log n)$, the total number of bits of certificates assigned to the special vertices of $G_{A,\bar{A}}$ is $o(n^{2})$, and thus the number of possible certificate assignments to the special vertices of $G_{A,\bar{A}}$ is $2^{o(n^{2})}$. Since there are $2^{(2^{k})^{2}}=2^{\Omega(n^{2})}$ choices for the set $A$, it follows from the pigeonhole principle that there exist two distinct sets $A,B$ such that the certificates of $G_{A,\bar{A}}$ and $G_{B,\bar{B}}$ agree on their special vertices. As $G_{A,\bar{A}}$ and $G_{B,\bar{B}}$ are both non-3-colorable, in any local certification scheme for non-3-colorability in these two graphs, each vertex accepts the instance upon reading its local view (i.e., the certificates of its closed neighborhood).
Note that at least one of $A\cap\bar{B}$ and $\bar{A}\cap B$ is non-empty, say $A\cap\bar{B}\neq\emptyset$ by symmetry. Consider the graph $G_{A,\bar{B}}$, and assign to the vertices of $V_{A}$ their certificates in $G_{A,\bar{A}}$, and to the vertices of $V_{\bar{B}}$ their certificate in $G_{B,\bar{B}}$. Observe that the local view of each vertex of $G_{A,\bar{B}}$ is the same as in $G_{A,\bar{A}}$ or $G_{B,\bar{B}}$, and thus each vertex accepts the instance. By definition of a proof labeling scheme, this implies that $G_{A,\bar{B}}$ is non-3-colorable. On the other hand, by (4) above, $G_{A,\bar{B}}$ is 3-colorable (because $A\cap\bar{B}\neq\emptyset$). This contradiction shows that non-3-colorability cannot be certified locally with certificates of size $o(n^{2}/\log n)$, as desired.
@@
For any integer $k$, and any two subsets $A\subseteq I\times I$ and $B\subseteq I\times I$ (where $I=\{1,\ldots,2^{k}\}$), we consider the graph $G_{A,B}$ of Göös and Suomela [15] introduced above, and we let $H_{A,B}=f(G_{A,B})$. Note that $H_{A,B}$ contains $n=\Theta(2^{2k})$ vertices, and moreover:
@@
    Each edge between $V_{A}$ and $V_{B}$ in $G_{A,B}$, connects a vertex of $S_{A}$ to a vertex of $S_{B}$, and the subgraph of $G_{A,B}$ induced by the special vertices $S_{A}\cup S_{B}$ is independent of the choice of $A$ and $B$ (in particular the size of $S_{A}$ and $S_{B}$ is independent of $A$ and $B$).
@@
The fact that the sets $S_{A}$ and $S_{B}$ still have size $O(\log n)$ in $H_{A,B}$ follows from item (3) in the properties of $G_{A,B}$ (the fact that there are only $O(\log n)$ edges with one endpoint in $S_{A}$ and one endpoint in $S_{B}$).
As above, if non-3-colorability can be certified locally with certificates of size $o(n/\log n)$ in graphs of maximum degree 4, the total number of bits of certificates assigned to the special vertices of $H_{A,\bar{A}}$ is $o(n)$. Since there are $2^{(2^{k})^{2}}=2^{\Omega(n)}$ choices for the sets $A$, it follows from the pigeonhole principle that there exist two distinct sets $A,B$ such that the certificates of $H_{A,\bar{A}}$ and $H_{B,\bar{B}}$ agree on their special vertices. We then obtain a contradiction using the same argument as above. ∎
@@
We now consider $P^{\prime}_{A,B}=f(P_{A,B})$ (the graph obtained from $P_{A,B}$ by applying the maximum degree 4 reduction of the proof of Theorem 3.2). This graph is a planar graph with maximum degree 4, is 3-colorable if and only if $A\cap B\neq\emptyset$, has $n=O(2^{4k})$ vertices, but only contains $O(\log^{2}n)$ special vertices. There are $2^{(2^{k})^{2}}=2^{2^{2k}}=2^{\Omega(\sqrt{n})}$ choices for the set $A$, so if we can locally certify non-3-colorability with certificates of size $o(n^{1/2}/\log^{2}n)$ then the total number of bits of certificates assigned to the special vertices of $P^{\prime}_{A,\bar{A}}$ is $o(n^{1/2})$ and thus there are two distinct sets $A,B$ such that the certificates of $P^{\prime}_{A,\bar{A}}$ and $P^{\prime}_{B,\bar{B}}$ agree on their special vertices. We then obtain a contradiction using the same argument as above. ∎